Analyze the annual reports of Flour Mills of Fiji (2020), RB Patel Ltd (2020) and Paradise Beverages Ltd (2019) from South Pacific Stock Exchange (SPSE) and answer the following questions. 

 

1.     Calculate the share value using Dividend Discount Model, assuming dividend will grow 3% for Flour Mills of Fiji, 6% for RB Patel Ltd and 9% for Paradise Beverages constantly for the companies on their respective base year. 

 

2.     Calculate the Earnings Per Share (EPS) and the Market Capitalization for each of the companies. 

 

3.     Comment on the share value of the companies given the Dividend Discount Model (Question 1) and EPS (Question 2).  


1
Expert's answer
2021-04-26T21:16:18-0400

(1)

Let the discount rate be 10, the dividends are 5,6,7, respectively, then we will find the price.

"P=\frac{D}{r-g}"

"P_1=\frac{5}{0.1-0.03}=71.43"


"P_2=\frac{6}{0.1-0.06}=150"


"P_1=\frac{7}{0.1-0.09}=700"


(2)

Let the net profit be 7000,8000,9000, respectively, and the number of shares 70,80,90, then the profit per share will be

"EPS=\frac{net profit}{s}"


"EPS_1=\frac{7000}{70}=100"


"EPS_2=\frac{8000}{80}=100"


"EPS_2=\frac{9000}{90}=100"


Market cap =share price × share outstanding

Market cap 1"=71.43\u00d770=5000.1"

Market cap 2"=150\u00d780=12000"

Market cap 3"=700\u00d790=63000"


(3)

The higher the dividend growth, the higher the share price.